OpenStudy (anonymous):

Someone help please! I need help making a table that shows the height from time t=0, here is the height versus time t: h = 70 + 150 t - 16 t^2

OpenStudy (whpalmer4):

Okay, so your formula is \[h(t) = 70+150t-16t^2\]You need to make a table showing values of \(t\) and the corresponding values of \(h(t)\) starting at \(t = 0\) and going probably until \(h(t) \le 0\) \[ \begin{array}{cc} t & h(t)\\\hline 0 & 70+150(0)-16(0)^2 = 70\\ 1 & 70+150(1)-16(1)^2 = \\ 2 & \text{etc.}\\ 3 & \\ 4 & \\ 5 & \\ 6 & \\ 7 & \\ 8 & \\ 9 & \\ 10 & \\ \end{array} \]